# ”Security for Everyone” in Finite Blocklength IRS-aided Systems With Perfect and Imperfect CSI

### v2

@@ III-3 Complexity Analysis
The developed Algorithm 1 is designed to tackle problem $(\mathcal{P}1)$ by decoupling the beamforming vector and the IRS’s PREs within the objective function. The resulting problem $(\mathcal{P}1)$ is an SDP problem that can be solved by the interior point method[30]. The algorithm’s complexity can be estimated by its worst-case runtime and the number of the decision variables [31]. Thus, in Algorithm 1, the computational complexity of obtaining $\boldsymbol{w}$ given $\boldsymbol{\theta}$ is $\mathcal{O}(M^{3})$, and the computational complexity of obtaining $\boldsymbol{\theta}$ given $\boldsymbol{w}$ is $\mathcal{O}(N^{3}(N+1))$.
IV MAXMIN SR under Imperfect CSI with FBR systems
To deal with the imperfect CSI from the IRS to the users and Eve, we adopt the channel modeling in equation (3) where only partial/imperfect CSI $\widehat{\boldsymbol{u}}_{i}$ is available. First, we can cast the minimum SR maximization problem as follows:
$(\mathcal{P}2):$ $\underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \underset{k\in\mathcal{K}}{\min}\leavevmode\nobreak\ \mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}),$ (35a)
@@ IV MAXMIN SR under Imperfect CSI with FBR systems
$\|\Delta\boldsymbol{u}_{i}\|_{2}\leq\xi_{i},i\in\{k,e\}.$ (35d)
Unlike the traditional LBR case under imperfect CSI, problem $(\mathcal{P}2)$ is more challenging. First, this is because the additional FBR constraints that are captured by the non-convex dispersion factor expression (12) are embedded within the objective function. Moreover, due to the imperfect CSI as modeled in (3), the semi-infinite constraint (35d) further complicates the problem.
To solve problem $(\mathcal{P}2)$, we first introduce slack variables to decompose the coupling of the beamforming vector and the IRS’s PREs in the objective function to facilitate the AO method. Specifically, we first substitute (13) into (35a), and then we introduce slack variables $z$ as the FBR-SR’s lower bound, ${\varphi}_{k}$ represents the minimum users’ rate, $\mu_{k_{e}}$ represents the maximum eavesdropping rate of Eve, and $\tilde{\varphi}_{i},i\in\{k,e\}$ represents the maximum dispersion factor for the user-$k$ and Eve, respectively. The problem $(\mathcal{P}2)$ can be recast as:
$(\mathcal{P}2.1):$ $\leavevmode\nobreak\ \underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ z,$ (36a)
s.t. $z\leq{\varphi}_{k}-\mu_{k_{e}}-\tilde{\varphi}_{k}-\tilde{\varphi}_{k_{e}},\forall k,$ (36b)
${\varphi}_{k}\leq\widehat{C}_{k},\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\xi_{k},\forall k,$ (36c)
$\mu_{k_{e}}\geq\widehat{C}_{e},\forall\|\Delta\boldsymbol{u}_{e}\|_{2}\leq\xi_{e},\forall k,$ (36d)
$\tilde{\varphi}_{k}\geq\xi_{k}\sqrt{\widehat{V}_{k}},\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\xi_{k},\forall k,$ (36e)
$\tilde{\varphi}_{k_{e}}\geq\xi_{e}\sqrt{\widehat{V}_{e}},\forall\|\Delta\boldsymbol{u}_{e}\|_{2}\leq\xi_{e},\forall k,$ (36f)
@@
One can notice that the constraint (35d) in problem $(\mathcal{P}2)$ has been captured by the constraints (36c), (36d), (36e), and (36f) in problem $(\mathcal{P}2.1)$ (referring to the CSI factor in the channel definition (4)). Next, we employ the SCA technique and the $\mathcal{S}$-procedure to transform the semi-infinite non-convex constraints (36c), (36d), (36e), and (36f) to finite LMIs. Then, we leverage a PCCP algorithm to tackle the UMC (35c) [32].
@@ IV-1 Sub-Problem for Optimizing the Beamforming Vectors
$2^{\varphi_{k}}-1\leq\frac{\left|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}}{\left\|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{k}},\forall k.$ (37)
By treating the interference plus noise signal as an auxiliary function $\boldsymbol{\beta}=[\beta_{1},\dots,\beta_{K}]$, (37) can be expressed as:
$\left|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\geq(2^{\varphi_{k}}-1)\beta_{k},\forall k,$ (38a)
$\left\|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{k}\leq\beta_{k},\forall k.$ (38b)
To circumvent the non-convex semi-infinite inequalities in (38a), we replace the left-hand side of (38a) with their lower bounds using the following Lemma.
@@
At iteration $(\iota)$, let $\boldsymbol{w}^{(\iota)}$ and $\boldsymbol{\theta}^{(\iota)}$ be the optimal solution, then at the point $(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ we can express the linear lower bound of (38a) as:
$\left|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\triangleq\boldsymbol{u}_{k}\mathbf{X}_{k}\boldsymbol{u}_{k}^{H},$ (39)
where $\mathbf{X}_{k}$, is given in (40).
@@
Next, by substituting (39) in (38a), and using (3a) and Lemma 1, the inequality in (38a) is reformulated as in (41), where $d_{k}\triangleq\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k}\widehat{\boldsymbol{u}}_{k}^{H}$.
To address the uncertainty of $\{\Delta\boldsymbol{u}_{k}\}$ in (41), we leverage the $\mathcal{S}$-procedure [16].
@@
$\begin{bmatrix}\mathbf{U}_{0}&\mathbf{u}_{0}\\ \mathbf{u}_{0}^{H}&\text{u}_{0}\end{bmatrix}-\sum\limits_{i=0}^{Q}\mathfrak{n}_{i}\begin{bmatrix}\mathbf{U}_{i}&\mathbf{u}_{i}\\ \mathbf{u}_{i}^{H}&\text{u}_{i}\end{bmatrix}\succeq 0.$ (43)
@@
$\begin{bmatrix}\varpi_{k}\mathbf{I}_{M}+\mathbf{X}_{k}&(\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})^{H}\\ (\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})&d_{k}-(2^{\varphi_{k}}-1)\beta_{k}-\eta_{k}\Omega_{k}^{2}\end{bmatrix}\succeq 0,\forall k,$ (44)
where $\boldsymbol{\eta}\triangleq[\eta_{1},\dots,\eta_{K}]^{T}\geq 0$ are slack variables. Even though (38a) has been transformed into an LMI form, it is still non-convex due to the non-convexity nature of the term $2^{\varphi_{k}}\beta_{k}$. For that, we adopt the SCA technique [15] to convert the non-convex constraint (38a) to a convex approximation expression. Specifically, performing the first order Taylor approximation, the term $2^{\varphi_{k}}\beta_{k}$ can be upper bounded by:
$(\beta_{k}(2^{\varphi_{k}}))^{ub}\triangleq((\varphi_{k}-\varphi_{k}^{(\iota)})(\beta_{k}^{(\iota)})\ln 2+\beta_{k})2^{\varphi_{k}^{(\iota)}},$ (45)
where $\varphi_{k}^{(\iota)},\beta_{k}^{(\iota)}$ are the value of the variables $\varphi_{k},\beta_{k}$ at iteration $(\iota)$ in the SCA-based algorithm, respectively. Lastly, substituting (45) in (44), the LMIs in (44) can be expressed as:
$\begin{bmatrix}\varpi_{k}\mathbf{I}_{M}+\mathbf{X}_{k}&(\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})^{H}\\ (\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})&d_{k}-(\beta_{k}(2^{\varphi_{k}}))^{ub}+\beta_{k}-\varpi_{k}\Omega_{k}^{2}\end{bmatrix}\succeq 0,\forall k.$ (46)
@@
$\begin{bmatrix}\beta_{k}&\mathbf{t}_{k}^{H}\\ \mathbf{t}_{k}^{H}&\mathbf{I}_{K-1}\end{bmatrix}\succeq 0,\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\Omega_{k},\forall k,$ (48)
where $\mathbf{t}_{k}\triangleq\left((\widehat{\boldsymbol{u}}_{k}^{H}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\mathbf{W}_{-k}\right)^{H}$, and $\mathbf{W}_{-k}\triangleq[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{k-1},\boldsymbol{w}_{k+1},\dots,\boldsymbol{w}_{\mathcal{K}}]\in\mathbb{C}^{M\times\mathcal{K}-1}$.
@@
$\begin{bmatrix}B_{11,k}&\widehat{t}^{H}_{k}&\mathbf{0}_{1\times M}\\ \widehat{t}_{k}&\mathbf{I}_{K-1}&\Omega_{k}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{k}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})&\varkappa_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (51)
where $B_{11,k}\triangleq\beta_{k}-\sigma_{k}-\varkappa_{k}$, and $\widehat{t}_{k}\triangleq\left((\widehat{\boldsymbol{u}}_{k}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\mathbf{W}_{-k}\right)^{H}$.
@@
Next, we tackle (36d) by firstly substituting (10) in (36d), and then treat the interference plus noise signal in (36d) as an auxiliary function $\boldsymbol{\beta}_{e_{k}}\triangleq[\beta_{e_{1}},\dots,\beta_{e_{k}}]$. We hence can recast (36d) as
@@
$\left|(\boldsymbol{u}_{e}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\leq(2^{\mu_{k_{e}}}-1)\beta_{ke},\forall k,$ (52a)
@@
$\left\|(\boldsymbol{u}_{e}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{e}\geq\beta_{ke},\forall k.$ (52b)
@@
To tackle the uncertainties of $\{\Delta\boldsymbol{u}_{e}\}$ in the constraints (52a) and (52b), we use a similar approach as in (38b). Hence, the equivalent LMIs of (52a) and (52b) are, respectively, given by,
@@
$\begin{bmatrix}C_{e,k}&\widehat{c}_{ke}^{H}&\mathbf{0}_{1\times M}\\ \widehat{c}_{ke}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k})&\vartheta_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (53)
@@
$\begin{bmatrix}C_{11,k}&-\widehat{c}_{ke}^{H}&\mathbf{0}_{1\times M}\\ -\widehat{c}_{ke}&\mathbf{I}_{K-1}&-\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})^{H}\\ \mathbf{0}_{M\times 1}&-\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})&\varrho_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (54)
@@
where $\boldsymbol{\vartheta}\triangleq[\vartheta_{1},\dots,\vartheta_{K}]\geq 0$ and $\boldsymbol{\varrho}\triangleq[\varrho_{1},\dots,\varrho_{K}]\geq 0$ are a slack variables,
$C_{e,k}$ $\triangleq(\beta_{ke}(2^{\mu_{k_{e}}}))^{ub}-\beta_{ke}-\vartheta_{k},$
$(\beta_{ke}(2^{\mu_{k_{e}}}))^{ub}$ $\triangleq((\mu_{k_{e}}-\mu_{k_{e}}^{(\iota)})(\beta_{ke}^{(\iota)})\ln 2+\beta_{ke})2^{\mu_{k_{e}}^{(\iota)}},$
$\widehat{c}_{ke}$ $\triangleq((\widehat{\boldsymbol{u}}_{e}^{H}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\mathbf{W}_{-k})^{H},$
$C_{11,k}$ $\triangleq\beta_{ke}-\sigma_{e}-\varrho_{k}.$

### v3

@@ III-3 Complexity Analysis
The developed Algorithm 1 is designed to tackle problem $(\mathcal{P}1)$ by decoupling the beamforming vector and the IRS’s PREs within the objective function. The resulting problem $(\mathcal{P}1)$ is an SDP that can be solved by the interior point method[31]. The algorithm’s complexity can be estimated by its worst-case runtime and the number of decision variables [32]. Thus, in Algorithm 1, the computational complexity of obtaining $\boldsymbol{w}$ given $\boldsymbol{\theta}$ is $\mathcal{O}(N^{3})$, and the computational complexity of obtaining $\boldsymbol{\theta}$ given $\boldsymbol{w}$ is $\mathcal{O}(M^{3}(M+1))$. One can notice that the algorithm complexity increases with $M$.
IV MAXMIN FBR-SR under Imperfect CSI
To deal with the imperfect CSI from the IRS to the users and Eve, we adopt the channel modeling in equation (3) where only partial/imperfect CSI $\widehat{\boldsymbol{l}}_{i}$ is available. First, we can cast the minimum SR maximization problem as follows:
$(\mathcal{P}2):$ $\underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \underset{k\in\mathcal{K}}{\min}\leavevmode\nobreak\ \widehat{\mathcal{S}}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}),$ (35a)
@@ IV MAXMIN FBR-SR under Imperfect CSI
$\|\Delta\boldsymbol{l}_{i}\|_{2}\leq\Omega_{i},i\in\{k,e\}.$ (35d)
$\tau_{k}\leq\tau_{\max},\leavevmode\nobreak\ \tau_{e}\leq\tau_{e_{\max}},\leavevmode\nobreak\ t_{t}\leq t_{\max},$ (35e)
where (35d) captures the channel estimation error. Unlike the traditional LBR case under imperfect CSI, problem $(\mathcal{P}2)$ is more challenging. First, this is because the additional FBR constraints, which are captured by the nonconvex dispersion factor expression (12), are embedded within the objective function. In addition, the imperfect CSI condition (35d), introduces uncertainty in the IRS-users/Eve’s channel, making the decoding error and consequently the inverse Q-function evaluation unreliable. This sensitivity significantly complicates the objective function, besides its nonlinearity and nonconvexity.
To solve problem $(\mathcal{P}2)$, we first introduce slack variables to decompose the coupling of the beamforming vector and the IRS’s PREs in the objective function to facilitate the AO method. Specifically, we first substitute (13) into (35a), and then we introduce slack variables $\mathcal{Z}$ as the FBR-SR’s lower bound, ${\beta}_{k}$ represents the minimum users’ rate, $\Upsilon_{k_{e}}$ represents the maximum eavesdropping rate of Eve, and $\tilde{\beta}_{i},i\in\{k,e\}$ represents the maximum dispersion factor for the user-$k$ and Eve, respectively. The problem $(\mathcal{P}2)$ can be recast as:
$(\mathcal{P}2.1):$ $\leavevmode\nobreak\ \underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \mathcal{Z},$ (36a)
s.t. $\mathcal{Z}\leq{\beta}_{k}-\Upsilon_{k_{e}}-\tilde{\beta}_{k}-\tilde{\beta}_{k_{e}},\forall k,$ (36b)
${\beta}_{k}\leq\widehat{C}_{k},\forall\|\Delta\boldsymbol{l}_{k}\|_{2}\leq\xi_{k},\forall k,$ (36c)
$\Upsilon_{k_{e}}\geq\widehat{C}_{e},\forall\|\Delta\boldsymbol{l}_{e}\|_{2}\leq\xi_{e},\forall k,$ (36d)
$\tilde{\beta}_{k}\geq\xi_{k}\sqrt{\widehat{V}_{k}},\forall\|\Delta\boldsymbol{l}_{k}\|_{2}\leq\xi_{k},\forall k,$ (36e)
$\tilde{\beta}_{k_{e}}\geq\xi_{e}\sqrt{\widehat{V}_{e}},\forall\|\Delta\boldsymbol{l}_{e}\|_{2}\leq\xi_{e},\forall k,$ (36f)
@@
Problem $\mathcal{P}2.1$ is still nonconvex due to the coupling between $\boldsymbol{w}$ and $\boldsymbol{\theta}$ within the constraints (36b), (36c), (36d), (36e), and (36f). In addition, the presence of the nonlinear square-root terms of the dispersion factors in (36e) and (36f) further increases the complexity of the problem. One can notice that the constraint (35d) in problem $(\mathcal{P}2)$ has been captured by the constraints (36c), (36d), (36e), and (36f) in problem $(\mathcal{P}2.1)$ (referring to the CSI factor in the channel definition (4)). Next, we linearize the square-root of the dispersion factors, and then we employ the SCA technique, the $\mathcal{S}$-procedure, and the first order Taylor approximation to transform the semi-infinite nonconvex constraints (36c), (36d), (36e), and (36f) to finite LMIs. Then, we leverage a PCCP algorithm to tackle the UMC (35c) [33].
@@ IV-1 Sub-Problem for Optimizing the Beamforming Vectors
$2^{\beta_{k}}-1\leq\dfrac{\left|(\boldsymbol{l}_{k}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}\right|^{2}}{\left\|(\boldsymbol{l}_{k}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{k}},$ (37)
where $\boldsymbol{W}_{-k}=[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{k-1},\boldsymbol{w}_{k+1},\dots,\boldsymbol{w}_{\mathcal{K}}]\in\mathbb{C}^{M\times\mathcal{K}-1}$. By treating the interference plus noise signal as an auxiliary function $\boldsymbol{\alpha}=[\alpha_{1},\dots,\alpha_{K}]$, (37) can be expressed as:
$\left|(\boldsymbol{l}_{k}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}\right|^{2}\geq(2^{\beta_{k}}-1)\alpha_{k},\forall k,$ (38a)
$\left\|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{k}\leq\alpha_{k},\forall k.$ (38b)
To circumvent the nonconvex semi-infinite inequalities in (38a), we replace the left-hand side of (38a) with their lower bounds using the following Lemma.
@@
At iteration $(n)$, let $\boldsymbol{w}^{(n)}$ and $\boldsymbol{\theta}^{(n)}$ be the optimal solution, then at the point $(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})$ we can express the linear lower bound of (38a) as:
$\left|(\boldsymbol{l}_{k}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}\right|^{2}\triangleq\boldsymbol{l}_{k}\mathbf{A}_{k}\boldsymbol{l}_{k}^{H},$ (39)
where $\mathbf{A}_{k}$, is given in (40).
@@
Next, by substituting (39) in (38a), and using (3a) and Lemma 1, the inequality in (38a) is reformulated as in (41), where $d_{k}\triangleq\widehat{\boldsymbol{l}}_{k}\mathbf{X}_{k}\widehat{\boldsymbol{l}}_{k}^{H}$.
To address the uncertainty of $\{\Delta\boldsymbol{l}_{k}\}$ in (41), we leverage the $\mathcal{S}$-procedure [14].
@@
$\begin{bmatrix}\mathbf{U}_{0}&\mathbf{u}_{0}\\ \mathbf{u}_{0}^{H}&\text{u}_{0}\end{bmatrix}-{\sum}_{i=0}^{Q}\mathfrak{n}_{i}\begin{bmatrix}\mathbf{U}_{i}&\mathbf{u}_{i}\\ \mathbf{u}_{i}^{H}&\text{u}_{i}\end{bmatrix}\succeq 0.$ (43)
@@
$\begin{bmatrix}\eta_{k}\mathbf{I}_{M}+\mathbf{A}_{k}&(\widehat{\boldsymbol{l}}_{k}\mathbf{A}_{k})^{H}\\ (\widehat{\boldsymbol{l}}_{k}\mathbf{A}_{k})&d_{k}-(2^{\beta_{k}}-1)\alpha_{k}-\eta_{k}\Omega_{k}^{2}\end{bmatrix}\succeq 0,\forall k,$ (44)
where $\boldsymbol{\eta}\triangleq[\eta_{1},\dots,\eta_{\mathcal{K}}]^{T}\geq 0$ are slack variables. Even though (38a) has been transformed into an LMI form, it is still nonconvex due to the nonconvexity nature of the term $2^{\beta_{k}}\alpha_{k}$. For that, we adopt the SCA technique [13] to convert the nonconvex constraint (38a) to a convex approximation expression. Specifically, performing the first order Taylor approximation, the term $2^{\beta_{k}}\alpha_{k}$ can be upper bounded by:
$(\alpha_{k}(2^{\beta_{k}}))^{ub}\triangleq\left((\beta_{k}-\beta_{k}^{(n)})(\alpha_{k}^{(n)})\ln 2+\alpha_{k}\right)2^{\beta_{k}^{(n)}},$ (45)
where $\beta_{k}^{(n)},\alpha_{k}^{(n)}$ are the value of the variables $\beta_{k},\alpha_{k}$ at iteration $(n)$ in the SCA-based algorithm, respectively. Lastly, substituting (45) in (44), the LMIs in (44) can be expressed as:
$\begin{bmatrix}\eta_{k}\mathbf{I}_{M}+\mathbf{A}_{k}&(\widehat{\boldsymbol{l}}_{k}\mathbf{A}_{k})^{H}\\ (\widehat{\boldsymbol{l}}_{k}\mathbf{A}_{k})&d_{k}-(\alpha_{k}(2^{\beta_{k}}))^{ub}+\alpha_{k}-\eta_{k}\Omega_{k}^{2}\end{bmatrix}\succeq 0,\forall k.$ (46)
@@
$\begin{bmatrix}\alpha_{k}&\mathbf{t}_{k}^{H}\\ \mathbf{t}_{k}^{H}&\mathbf{I}_{K-1}\end{bmatrix}\succeq 0,\forall\|\Delta\boldsymbol{l}_{k}\|_{2}\leq\Omega_{k},\forall k,$ (48)
where $\mathbf{t}_{k}\triangleq\left((\widehat{\boldsymbol{l}}_{k}^{H}\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\mathbf{W}_{-k}\right)^{H}$.
@@
$\begin{bmatrix}Q_{11,k}&\widehat{t}^{H}_{k}&\mathbf{0}_{1\times M}\\ \widehat{t}_{k}&\mathbf{I}_{K-1}&\Omega_{k}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}\boldsymbol{W}_{-k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{k}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}\boldsymbol{W}_{-k})&\varkappa_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (51)
@@
where $Q_{11,k}\triangleq\alpha_{k}-\sigma_{k}-\varkappa_{k}$, and $\widehat{t}_{k}\triangleq\left((\widehat{\boldsymbol{l}}_{k}\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\mathbf{W}_{-k}\right)^{H}$. Next, we tackle (36d) by firstly substituting (10) in (36d), and then treat the interference plus noise signal in (36d) as an auxiliary function $\boldsymbol{\alpha}_{e_{k}}\triangleq[\alpha_{e_{1}},\dots,\alpha_{e_{k}}]$. We hence can recast (36d) as
@@
$\left|(\boldsymbol{l}_{e}\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}\right|^{2}\leq(2^{\Upsilon_{k_{e}}}-1)\alpha_{ke},\forall k,$ (52a)
@@
$\left\|(\boldsymbol{l}_{e}\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{e}\geq\alpha_{ke},\forall k.$ (52b)
@@
To tackle the uncertainties of $\{\Delta\boldsymbol{l}_{e}\}$ in the constraints (52a) and (52b), we use a similar approach as in (38b). Hence, the equivalent LMIs of (52a) and (52b) are, respectively, given by,
@@
$\begin{bmatrix}C_{e,k}&\widehat{b}_{ke}^{H}&\mathbf{0}_{1\times M}\\ \widehat{b}_{ke}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}\boldsymbol{w}_{k})&\vartheta_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (53)
@@
$\begin{bmatrix}B_{11,k}&-\widehat{b}_{ke}^{H}&\mathbf{0}_{1\times M}\\ -\widehat{b}_{ke}&\mathbf{I}_{K-1}&-\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}\boldsymbol{W}_{-k})^{H}\\ \mathbf{0}_{M\times 1}&-\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}\boldsymbol{W}_{-k})&\varrho_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (54)
@@
where $\boldsymbol{\vartheta}\triangleq[\vartheta_{1},\dots,\vartheta_{K}]\geq 0$ and $\boldsymbol{\varrho}\triangleq[\varrho_{1},\dots,\varrho_{K}]\geq 0$ are a slack variables, and,
$C_{e,k}$ $\triangleq(\alpha_{ke}(2^{\Upsilon_{k_{e}}}))^{ub}-\alpha_{ke}-\vartheta_{k},$
$(\alpha_{ke}(2^{\Upsilon_{k_{e}}}))^{ub}$ $\triangleq((\Upsilon_{k_{e}}-\Upsilon_{k_{e}}^{(n)})(\alpha_{ke}^{(n)})\ln 2+\alpha_{ke})2^{\Upsilon_{k_{e}}^{(n)}},$
$\widehat{b}_{ke}$ $\triangleq((\widehat{\boldsymbol{l}}_{e}^{H}\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\mathbf{W}_{-k})^{H},$
$B_{11,k}$ $\triangleq\alpha_{ke}-\sigma_{e}-\varrho_{k}.$